OpenStudy (anonymous):

Find all the complex roots of x^3+x=0

OpenStudy (anonymous):

\(x^3+x=x(x^2+1)\) by pulling out the common factor

OpenStudy (anonymous):

can you solve \(x^2+1=0\)?

OpenStudy (anonymous):

no im bad at algebra

OpenStudy (misty1212):

lol

OpenStudy (misty1212):

\[x^2+1=0\\ x^2=-1\\ x=\pm\sqrt{-1}\] which you can write as \[x=\pm i\] if you like

OpenStudy (anonymous):

so is that the answer?

OpenStudy (misty1212):

yeah

OpenStudy (anonymous):

oh, thank you!!

OpenStudy (misty1212):

is it multiple choice or do you have to write it?